Chromium metal (Cr) can be prepared by reacting the
chromium(III) oxide (Cr2O3) with aluminum (Al) according this reaction:
Cr2O3 (s) + 2Al (s)  2Cr (s) + Al2O3 (s)
If 100 g of Cr2O3 and 100 g of Al are reacted
What is a limiting reactant (L.R)?
1
Expert's answer
2020-11-26T05:21:51-0500

To find the limiting reagent, look for mole of each reagent that will react, the reagent with the smaller mole per coefficient ratio is the limiting reagent

  • Reagent Cr2O3 Al
  • Molar mass 151.99g 26.982g
  • Reacting mass 100g 100g
  • Mole. 100/151.99 = 0.658. 100/26.982 = 1.853
  • Mole per coefficient ratio. 0.658 3.706/2 = 1.853

Therefore, Cr2O3 is the limiting reagent